# Depression Detection on Social Media with Large Language Models

### v1

@@ 4.2.3 Hyperparameter Study
In our method design, we introduce two hyperparameters: $k$, which controls the proportion of text retained with a high possibility of depression symptoms, and $m$, which controls the proportion of text retained with high emotional intensity. In this section, we explore the impact of these two parameters on model performance. In our experiments, $k$ and $m$ were varied from 0 to 100 in increments of 10, corresponding to retaining 0% to 100% of the texts. The experimental results are shown in Figure 3.
@@
  • Impact of $k$. The parameter $k$ determines the proportion of text retained with a high risk of depression. As $k$ increases, the AUPRC monotonically increases. This is because the more text is annotated, the more information is provided. Annotating 20% of the text with the LLM, compared to not annotating at all, significantly improves AUPRC, and it is already close to the performance when all texts are annotated. This demonstrates the rationality of our design for efficient implementation.
  • Impact of $m$. The parameter $m$ determines the proportion of text retained with high emotional intensity. As $m$ increases, the AUPRC initially rises and then falls. The selection of $m$ needs to balance between sufficient filtering effect and not losing too much information. However, regardless of the value of $m$, constructing a mood course representation always benefits performance. This validates the rationality of our method design.
4.2.4 Case Study
In our system, we annotate symptoms exhibited in users’ posts, summarized the users’ mood course, and based on these, provide explanations for the system’s final decision output. An example is presented in Figure 4 to aid in better understanding the operation of our system.
Figure 4: A case study of DORIS’s output.
As seen, during the diagnostic criteria feature construction stage, our method accurately identifies depression-related symptoms in posts. In the mood course representation construction stage, our approach selects posts with high emotional density, thereby generating a mood course description. By combining identified symptoms of depression and the user’s mood course, the system can generate a rational explanation for its judgment.
5 Conclusion and Future Work
In this work, we introduce a novel depression detection system named DORIS, which incorporates clinical practices in depression diagnosis, including the widely used diagnostic criteria DSM-5, and analyses of mood course, achieving high accuracy and interpretability in depression detection. Extensive experimental results demonstrate the effectiveness of our method, as well as the rationality of each design component in DORIS.
In the future, we plan to deploy our system in real-world settings to collaborate with human experts in clinical diagnosis. This will enable us to identify areas for improvement in our system and continue its iterative design.
Limitations
  • Our experiment is based on a single dataset. Large-scale datasets with expert annotations are challenging to acquire in the field of depression detection, and we find only one dataset that meets our criteria, further refined based on our standards. The experimental results on this dataset have demonstrated the effectiveness of our method. However, we believe that experiments on more datasets could further validate our approach.
  • Our focus is solely on depression. Mental health analysis and mental illness detection are crucial for human well-being. Our work concentrates on only one of these tasks: depression detection. While depression is a major mental illness, extending our method to detect other mental health illnesses, such as anxiety and bipolar disorder, would be more meaningful. Specifically, our explicit modeling of the mood course could be significantly beneficial for accurate and interpretable detection of bipolar disorder. In the future, we plan to extend our method to a broader range of mental health analysis and mental illness detection tasks.
Ethical Considerations
The texts presented in our paper include content with depressive tendencies, which may cause discomfort. The dataset used is a public dataset, and we have strictly adhered to the guidelines provided by the dataset providers. To protect user privacy and prevent misuse, all examples shown in the paper have been moderately obfuscated bruckman2002studying . In using the API services provided by OpenAI, we have complied with OpenAI’s terms and policies.
Appendix A Discussions on Our Utilized Dataset
Unlike other machine learning tasks with large-scale, high-quality datasets with golden labels, acquiring well-labeled, large-scale datasets for depression detection is challenging due to the sensitivity and privacy concerns associated with mental illnesses. Some datasets utilize self-reported medical diagnoses as the criterion for identifying users with depression. The advantage of this approach is its relatively high reliability and ease of data collection without needing expert involvement. However, this leads to datasets comprising users who are already medically diagnosed with depression and willing to discuss it online. However, the most valuable aspect of depression detection task is assisting human experts in identifying potentially undiagnosed users on the internet and intervening when appropriate. Moreover, posts from self-disclosing users often contain explicit mentions of depression or medication discussions, making identification relatively easy. We want our utilized dataset to include users who have not self-reported depression but are highly likely to be depressed, demonstrating the effectiveness of our method in such scenarios. Additionally, the ratio of depressed to normal users in most previous datasets [25, 49, 40] is around 1:4 or even 1:1, deviating far from the actual prevalence of about 5% in the population. To simulate a real-world application scenario, our dataset has a ratio of 1:19 depressed to normal users. Experiments on such a dataset can more adequately validate the practical performance of various methods.
Given the scarcity of large-scale open-source datasets in depression detection, finding a dataset that meets our criteria was challenging. Fortunately, we discovered a large-scale, open-source dataset, SWDD [7], supporting our experiments. This high-quality dataset has been carefully annotated by experts, including users who self-disclosed a medical diagnosis of depression and those identified by experts as highly likely to be depressed based on medical criteria, and without self-disclosure. We retain these non-self-disclosing users and, after further expert screening, keep 1000 depressed users and 19000 normal users. Following the findings of [16], that utilizing posts from the six months preceding a medical diagnosis, rather than the entire history, yields a higher accuracy in predicting clinically diagnosed depression; and due to the token window constraint of LLMs, we retain only the posts posted within six months prior to each user’s last post. The dataset statistics are shown in Table 2. Training, validation, and test sets were divided in a 7:1:2 ratio, as recommended by [40]. The original language of this dataset is Chinese. In our paper, to make our work accessible to a broader audience, all texts from the dataset are translated into English when presented.
Appendix B Symptom Templates
@@ F. Fatigue
Appendix C Emotion Templates
@@
I am sad, sorrowful, melancholic, in pain, lost, depressed, pessimistic, tearful, grieving, mournful, depressed, suicidal, heartbroken, devastated, upset, crying, deeply saddened, disconsolate, dejected, lamenting, desolate, gloomy, mournful, weeping bitterly, desperate, heartbroken, indignant.

### v2

@@ Appendix D Hyperparameter Study (RQ4)
In our method design, we introduce four hyperparameters: $k$ (proportion of text for LLM annotation), $m$ (proportion of emotionally intensive text), and $\alpha$ and $\beta$ (ratio between LLM analysis and original posts in mood course representation). The experimental results are shown in Figure 4.
@@
Impact of $k$. As $k$ increases, the AUPRC monotonically increases. Annotating 20% of the text with the LLM significantly improves AUPRC and is close to the performance when all texts are annotated. This demonstrates the rationality of our design for efficient implementation.
Impact of $m$. As $m$ increases, the AUPRC initially rises and then falls, suggesting a trade-off between capturing emotionally intensive posts and retaining sufficient information.
Impact of $\alpha$ and $\beta$. When $\alpha$$\beta$ ratio is 2:3, the system achieves optimal performance.
Appendix E Practical Design Considerations
Our research is motivated by the real-world need to address the global shortage of mental health services. Literature indicates an urgent need for effective, automated tools to assist with large-scale, low-cost early screening, complementing the efforts of human professionals Whitton et al. (2021); Balcombe and De Leo (2021). The design philosophy of DORIS is guided by deployability and considers several practical constraints:
Cost and Efficiency.
We designed an efficient filtering mechanism (Section 2.2.2) to significantly reduce expensive LLM API calls. As shown in the hyperparameter study (Appendix I), annotating only 20% of the text achieves performance close to full annotation. This is a critical consideration for large-scale deployment.
Privacy and Flexibility.
In addition to using APIs, we validated our framework’s effectiveness with locally deployed open-source LLMs (Appendix J). This directly addresses the core concerns of data privacy and security in real clinical settings and demonstrates deployment flexibility.
Explainability.
For medical applications, we view explainability not just as an academic pursuit but as a core prerequisite for gaining trust from clinicians and driving adoption. The hybrid design ensures that the system provides concrete evidence aligned with clinical practice.
Appendix F Prompt Instruction
The following prompt is used to instruct the LLM to identify depression symptoms within a given post:
Assuming you are a psychiatrist specializing in depression. Given [text], please determine if this message includes any of the following states of the author: A. Depressive mood B. Loss of interest/pleasure … I. Thoughts of suicide. If present, answer in the format of enclosed letters separated by commas, for example, (A, B, C). If none are present, respond with None.
The following prompt is used to instruct an LLM to synthesize a description of user $u$’s mood course, $T^{MC}$:
"As a consulting psychiatrist, please conduct a longitudinal mood course analysis based on the following temporal sequence of personal expressions. For each entry, evaluate affect, emotional valence, and severity of mood states. Synthesize these observations into a clinical summary of mood progression, noting any patterns of persistence, fluctuation, or changes over time: Time: $t_{1}$, Post: $p_{1}$, Time: $t_{2}$, Post: $p_{2}$, …"
The following prompt is used to instruct the LLM to generate the explanation text, $T^{Exp}$:
"Assuming you are a psychiatrist specializing in depression. Here is a user’s mood course: $T^{MC}$; below are posts from this user displaying symptoms of depression and the types of symptoms exhibited: …; this user has been determined by an automated depression detection system to be depressed/normal. Please consider the user’s mood course and posts to generate an explanation for this judgment. Your explanation should be grounded in concrete evidence."
Appendix G Symptom Templates
During the Diagnostic Criteria Feature Extraction phase, with the assistance of clinical psychologist and psychiatrists, we designed symptom templates to filter text with a high likelihood of depression through matching. Specifically, the templates we designed are as follows:
@@ F. Fatigue
Appendix H Emotion Templates
During the Mood Course Representation Construction phase, with the assistance of clinical psychologist and psychiatrists, we designed emotion templates. The purpose of emotion templates is to filter text with a high emotional intensity to generate the user’s mood course. Specifically, the templates we designed are as follows:
@@ Appendix H Emotion Templates
I am sad, sorrowful, melancholic, in pain, lost, pessimistic, tearful, grieving, mournful, depressed, suicidal, heartbroken, devastated, upset, crying, deeply saddened, disconsolate, dejected, lamenting, desolate, gloomy, weeping bitterly, desperate, heartbroken, indignant.